1. Find the general solution for each of the following differential equations (10 points each): y - 2y - 3y = 32 y - y - 2
Answer #1

Find the complemetary functions and particular integral for each of the following.

a Answert General solution for each of the following i y - 24 - 37 = 3 e 2x For Yeo yu-ay - 3y=0 Auxiliary Equations (CP-2р = 1 (-2141-) (- 25. 4-) 1 (D+1)(-2) - Бр-2 1 (- 2x +4х2) - 2 - D)) Чr = using Bonchaup epicy - - 2x +4х2) 1 + t 2 ч, г-(-a+For y = I (1263* + 12 -29) (D-2)(0+3) 12 e 32 t 12 e 2x (-2-2) (-2+3) (3-2)(3+3) 12e + -2x 20 3 x 3 e 2x 12 e -4 m² 2m-3 =0 YJoyet Up yo A ePt Be3a-32*1*4** ***) (e) y + 2y ty: 2 ex y = Ye + Yp For Y,= (b²+ 2D + 3y=0 Auxiliary equation : (m+1²=0 Jc